The expected return on the market is 10 percent, the risk-free rate is 4.3 percent, and the tax rate is 45 percent. Fairfax Paint has 1,400,000 common shares outstanding that are priced at 35 dollars per share and have an expected return of 19.51 percent and an expected real return of 17.25 percent. Last year, Fairfax Paint common stock had a return of 21.13 percent. The company also has 600,000 shares of preferred stock outstanding that are priced at 15 dollars per share and have an expected return of 12.12 percent and an expected real return of 10 percent. Last year, Fairfax Paint preferred stock had a return of 13.2 percent. Finally, the company has 10,000 bonds outstanding with a coupon rate of 9.79 percent, yield-to-maturity of 5.15 percent, current yield of 8.9 percent, face value of 1000 dollars, and price of 1,100 dollars. What is the weighted average cost of capital for Fairfax Paint? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

Weighted average cost of capital=(weight of debt*after tax cost of debt)+(weight of equity*cost of equity)+(weight of preferred stock*cost of preferred stock)

Cost of equity=shareholders expected rate of return=19.51%

Cost of preferred stock=expected return=12.12%

Before tax cost of debt=yield to maturity=5.15%

After tax cost of debt=before tax cost of debt*(1-tax rate)=5.15%*(1-45%)=2.83%

Market value of equity=shares outstanding*stock price per share=1400000*35=49000000

Market value of preferred stock=number of shares*preferred stock price=9000000

Market value of debt=number of bonds*price of each bond=10,000*1100=11,000,000

Total value=49000000+9000000+11000000=69000000

Weight of equity=market value of equity/total value=49000000/69000000=71.01%

Weight of preferred stock=9000000/69000000=13.04%

Weight of debt=11000000/69000000=15.94%

Weighted average cost of capital=(71.01%*19.51%)+(13.04%*12.12%)+(15.94%*2.83%)

=15.89%